# Modeling isotropic polyconvex hyperelasticity by neural networks – sufficient and necessary criteria for compressible and incompressible materials

### 28 Mar 2026 v1

@@ 3.3.1 The compressible case
The proof is provided in A. This criterion bears similarities to Ball’s well-known criterion for polyconvexity. In recognition of Ball’s work, the energy function is therefore denoted by $\Psi_{\textnormal{Ball}}$. However, Ball’s criterion requires a stricter permutation invariance:
@@
$\Psi(\boldsymbol{F})=\psi_{(}\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{2}\lambda_{3},\lambda_{1}\lambda_{3},\lambda_{1}\lambda_{2},\lambda_{1}\lambda_{2}\lambda_{3}),$ (30)
@@
$\textnormal{(b) }\psi(x_{1},x_{2},x_{3},y_{1},y_{2},y_{3},\delta)\text{ is non-decreasing in each }x_{i},y_{j}.$ (32)
Then $\Psi$ is polyconvex on $U$.
@@
$\Psi_{\textnormal{Ball}}(\boldsymbol{B}\cdot\boldsymbol{\lambda},\boldsymbol{B}\cdot\boldsymbol{\check{\lambda}},\lambda_{1}\lambda_{2}\lambda_{3})=\Psi_{\textnormal{Ball}}(\boldsymbol{\lambda},\boldsymbol{\check{\lambda}},\lambda_{1}\lambda_{2}\lambda_{3})\quad\forall\,\boldsymbol{B}\in\textnormal{Perm}(3)$ (33)
since $\boldsymbol{\lambda}$ and $\boldsymbol{\check{\lambda}}$ as well as their permutations are connected, which can be derived from Theorem 1. Nevertheless, this criterion is still only sufficient but not necessary and rules out some frame-indifferent, isotropic polyconvex functions as numerical examples in Section 7 will show.
@@ 3.3.2 The incompressible case
The derivation of the above criterion for the incompressible case is also straightforward.
@@ 3.4.1 The compressible case
A criterion formulated in matrix invariants can be directly motivated from Ball’s criterion. Within this work, the matrix invariants ($I_{1}^{\boldsymbol{U}}=\lambda_{1}+\lambda_{2}+\lambda_{3}$, $I_{2}^{\boldsymbol{U}}=\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}+\lambda_{1}\lambda_{2},I_{3}^{\boldsymbol{U}}=J=\lambda_{1}\lambda_{2}\lambda_{3}$) of the right stretch tensor $\boldsymbol{U}=\sqrt{\boldsymbol{C}}=\sqrt{\boldsymbol{F^{T}}\cdot\boldsymbol{F}}$ are used; see, for example, [40] for more details. Let $\Psi$ be given by
@@
Certainly, the invariants of the right Cauchy–Green tensor $\boldsymbol{C}$ could also be used. However, it was already shown in [16] that this criterion is not able to approximate, e.g., energy functions that grow linearly with respect to the stretch. This results in stress plateaus, which are observed, e.g., for polymers and biological materials such as the periodontal ligament [37, 24].
@@ 3.4.2 The incompressible case
The criterion reduces to $I_{1}^{\boldsymbol{U}}=\lambda_{1}+\lambda_{2}+\lambda_{3}$ and $I_{2}^{\boldsymbol{U}}=\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}+\lambda_{1}\lambda_{2}$ for incompressibility. Let $\Psi^{\textnormal{inc}}$ be given by
@@ 4 Isotropic polyconvex neural networks for incompressibility
where $\boldsymbol{x}$ is the input vector, $\boldsymbol{z}_{i}$ are the layer activations ($\boldsymbol{z}_{0}\equiv 0,$ $\boldsymbol{W}_{0}^{(\boldsymbol{z})}\equiv 0$), $\theta=\{\boldsymbol{W}_{0:k-1}^{(\boldsymbol{x})},\boldsymbol{W}_{1:k-1}^{(\boldsymbol{z})},\boldsymbol{b}_{0:k-1}\}$ are the weights, $g_{i}$ are non-linear activation functions and $k$ is the number of layers. Given this architecture, $\mathcal{NN}$ is convex in $\boldsymbol{x}$ if all $\boldsymbol{W}_{1:k-1}^{(\boldsymbol{z})}$ are non-negative and all activation functions $g_{i}$ are convex and non-decreasing [4]. ICNNs are employed in this work because of their representational power. A universal approximation theorem for convex functions was proven by Huang et al. [20] and Chen et al. [12] when ReLU or Softplus activation functions are used (see also Lemma 6). Of course, other neural network architectures ensuring convexity in the input arguments could be employed, see Remark 1. However, universal approximation theorems would have to be developed for these architectures if this is even possible.
@@ 4.1 Convex Signed Singular Value Neural Networks
To approximate any frame-indifferent, isotropic polyconvex function under the incompressibility constraint the same CSSV-NN framwork as in [16], but reduced by the determinant, is proposed and hence labeled as inc-CSSV-NN. This corresponds to the polyconvexity criterion given by Corollary 2. The modified inc-CSSV-NN framework is summarized briefly below.
@@
See Tab. 1 for all combinations of $\boldsymbol{x}^{(j)}$. It should be emphasized that the same $\mathcal{NN}$ and therefore the same weights have to be used for every input permutation in order to guarantee $\Pi_{3}$-invariance. The overall inc-CSSV-NN architecture and all following neural network architectures are sketched in Fig. 3. A universal approximation theorem for the underlying inc-CSSV-NN is proven in Section 5. Such a universal approximation theorem cannot be derived for the following neural networks since they are only based on sufficient criteria. The first Piola–Kirchhoff stress tensor defined by the network potential then reads
@@
Kolmogorov–Arnold Networks (KANs) [28] currently emerge as promising alternatives to ICNNs. They were first applied to constitutive modeling by Abdolazizi et al. [2] and afterwards extended to account for polyconvexity in [42]. While the Kolmogorov–Arnold representation theorem provides a strong theoretical background, the approximation capabilities of KANs, especially under further constraints such as convexity, and limitations are yet to be explored more deeply and universal approximation theorems (for those specific cases) are to be developed. While this states an interesting, future challenge, it is beyond the scope of this work and not necessary for the present purpose since ICNNs already guarantee universal approximation. Furthermore, the required Perm(3)(3)-invariance could also be ensured via Deep Sets [48] to eventually save some computational time. However, it has not yet been investigated whether universal approximation can be achieved for the underlying constraints of $\Pi_{3}$-invariance and convexity. The proposed average technique in Eq. 38 is simple, effective, does not limit the network expressiveness and is computationally efficient since the network evaluations can be performed simultaneously. Nevertheless, Deep Sets are another noteworthy option for future studies.
@@ 4.2 Reduced Convex Signed Singular Value Neural Networks
4.3 Neural networks based on Ball’s criterion for polyconvexity formulated in principal stretches
The general framework for neural network models based on Ball’s criterion for polyconvexity (here referred to as Ball NN and inc-Ball NN in the incompressible case) is very similar to the one shown in the previous sections. The elementary polynomials of the principal stretches are the network input instead of the signed singular values. Then, only the respective invariance has to be changed and the non-decreasing property of Corollary 5 needs to be taken into account. The Perm(3)(3)-invariance is guaranteed by setting
@@ 4.3 Neural networks based on Ball’s criterion for polyconvexity formulated in principal stretches
See Tab. 3 for all combinations of $\boldsymbol{x}_{\lambda}^{(j)}$. To account for the additional non-decreasing property – in addition to $\boldsymbol{W}_{1:k-1}^{(\boldsymbol{z})}$ – also all $\boldsymbol{W}_{0:k-1}^{(\boldsymbol{x})}$ have to be non-negative. This is quite restrictive for the expressiveness of ICNNs and comes on top of the fact that Ball’s criterion is only sufficient for polyconvexity. However and for the same reasons as the reduced inc-/CSSV-NN, the model is computationally more efficient and less susceptible to local minima during optimization than the inc-/CSSV-NN but less efficient compared to the reduced inc-/CSSV-NN.
Table 3: Input permutations of the principal stretches for the neural network employed in Eq. (44) based on Ball’s criterion for polyconvexity.
@@ 4.4 Polyconvex neural networks based on matrix invariants
However, as for the neural network models based on Ball’s criterion, all $\boldsymbol{W}_{0:k-1}^{(\boldsymbol{x})}$ have to be non-negative to ensure the non-decreasing property. While this model, henceforth referred to as inc-UInvar NN, is the simplest and most efficient computationally, it inherently has the least expressive power.
@@ 6.1 Network architecture
CSSV-NN 7 Ball NN 7
inc-CSSV-NN 6 inc-Ball NN 6
Reduced CSSV-NN 3 UInvar NN 3
Reduced inc-CSSV-NN 2 inc-UInvar NN 2
@@ 6.2 Sobolev training
In constitutive modeling, it is now common practice to train the networks in Sobolev space by strain–stress tuples
@@
with the Froebenius norm $\|\bullet\|$. This formulation measures deviations in stress space, thus offering a natural physical interpretation. While alternative loss definitions are possible (e.g., based on strain–energy data or a mixture of strain–energy and strain–stress), training exclusively on stress data produced the most reliable outcomes and aligns well with the availability of experimental measurements. Note that the inc-CSSV-NN framework can also be applied to approximate polyconvex hulls in the same manner as described in [16].
@@ 7 Numerical examples
In the following, several numerical examples are conducted to analyze the expressiveness, accuracy, simplicity and efficiency of the proposed models. In a first step, the models are trained on synthetic data from four well-established, classical models. Then, they are trained on Treloar’s experimental data [43] to highlight their practical applicability. Finally, an explicitly constructed energy function is provided that cannot be approximated by neural networks based on Ball’s criterion for polyconvexity.
@@ 7.1 Classical models
leading to 52 data points (with uniform distribution within each load case). $F_{33}$ follows from incompressibility and the Lagrange multiplier $p$ from the constraint $P_{33}=0$. Validation and test data sets are omitted for the noise-free synthetic data base that is directly extracted from the ground truth reference. This does not limit any of the following statements. For each combination of material model, neural network model (inc-CSSV-NN, reduced inc-CSSV-NN, inc-Ball NN and inc-UInvar NN) and architecture (1–8), 30 instances with different random weight initializations were trained using SciPy’s minimize function, which performs a constrained L-BFGS-B optimization. The number of maximum iterations was set to 1000. However, nearly every optimization stopped earlier based on SciPy’s default criteria.
@@ 7.2 Treloar data
Again, the inc-CSSV-NN, the reduced inc-CSSV-NN and the inc-Ball NN fit the material behavior perfectly, while the inc-UInvar NN fails to do so, cf. Fig. 8. Accordingly, this is also evident in the mean squared errors shown in Tab. 7. It should be mentioned, that the first three neural network models perform magnitudes better than all 14 classical models investigated in [41] highlighting the superior expressiveness and accuracy of neural network models. Unfortunately, an exact error comparison is not possible because the stress components included in the error [41] are not explicitly accessible. Interestingly, the test set mean squared errors are lower than the training and validation errors. This can be justified by the data distribution: the randomly selected, few test data points lie all in low stretch and stress regimes, which inherently show lower squared errors. In these regimes even the inc-UInvar NN provides a really good approximation.
So far, no significant difference between the inc-CSSV-NN, the reduced inc-CSSV-NN and the inc-Ball NN regarding the accuracy and expressiveness can be identified. For this reason, a final example is examined to do so.
@@ 7.3 Mielke energy
and showed that it is polyconvex, despite being non-convex in $\boldsymbol{\nu}$ and, in particular, not monotone in any $\nu_{i}$. Especially due to the latter, it seems reasonable to assume that this energy cannot be approximated by neural networks that are based on Ball’s criterion for polyconvexity, which is investigated in this section. Since the energy (55) is not differentiable, a smooth, polyconvex approximation given by
@@ 7.3.1 Incompressible Mielke energy
This time only the inc-CSSV-NN and the reduced inc-CSSV-NN are able to reproduce the material model perfectly, cf. Fig. 9 and Tab. 6. Their lowest mean squared errors are 9.73e-05 for the inc-CSSV-NN and 6.63e-05 for the reduced version. In contrast, the lowest mean squared error achieved by the inc-Ball NNs (6.95e-03) is 71 times higher than the lowest inc-CSSV-NN error and 104 times higher than the lowest reduced inc-CSSV-NN error. The inc-Ball NN is, e.g., not able to predict the stress component $P_{11}$ for compressive strains in $11$-direction while stretched in $22$-direction, see Fig. 9 right, or decreasing stresses in the pure shear case ($\lambda_{2}=1$, Fig. 9 middle and right) accurately. This behavior is expected from the respective polyconvexity criterion. The inc-UInvar NN produces even worse results (lowest mean squared error 7.44e-02) and shows only a smooth, but far off approximation of the stresses (Fig. 9). The details of the material behavior prescribed by Eq. (57) are not captured.
Finally, it can be concluded from this example that the expressivity of the inc-Ball NN is indeed limited indicating that Ball’s polyconvexity criterion is not necessary in the incompressible case. In contrast, the reduced inc-CSSV-NN approximated all incompressible energies perfectly, what does not contradict the ?-marked relation in Fig. 2 regarding the necessity of the reduced criterion based on the elementary symmetric polynomials of the signed singular values in the incompressible case. However, further mathematical investigations are required and rigorous proofs of these statements remain an open challenge.

### 13 Apr 2026 v2

@@ 3.3.1 The compressible case
The proof is provided in Appendix A. This criterion bears similarities to Ball’s well-known criterion for polyconvexity. In recognition of Ball’s work, the energy function is therefore denoted by $\Psi_{\textnormal{Ball}}$. However, Ball’s original criterion requires a stricter permutation invariance:
@@
$\Psi(\boldsymbol{F})=\psi(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{2}\lambda_{3},\lambda_{1}\lambda_{3},\lambda_{1}\lambda_{2},\lambda_{1}\lambda_{2}\lambda_{3}),$ (29)
@@
$\textnormal{(b) }\psi(x_{1},x_{2},x_{3},y_{1},y_{2},y_{3},\delta)\text{ is non-decreasing in each }x_{i},y_{j},$ (31)
wherein $\boldsymbol{x}=(x_{1},x_{2},x_{3})$ and $\boldsymbol{y}=(y_{1},y_{2},y_{3})$ denote the first three and the fourth through sixth arguments of $\psi$, respectively. Then $\Psi$ is polyconvex on $U$.
@@
$\Psi_{\textnormal{Ball}}(\boldsymbol{B}\cdot\boldsymbol{\lambda},\boldsymbol{B}\cdot\boldsymbol{\check{\lambda}},\lambda_{1}\lambda_{2}\lambda_{3})=\Psi_{\textnormal{Ball}}(\boldsymbol{\lambda},\boldsymbol{\check{\lambda}},\lambda_{1}\lambda_{2}\lambda_{3})\quad\forall\,\boldsymbol{B}\in\textnormal{Perm}(3)$ $\text{ with }\boldsymbol{\lambda}=(\lambda_{1},\lambda_{2},\lambda_{3})$ (32)
$\text{ and }\boldsymbol{\check{\lambda}}=(\lambda_{2}\lambda_{3},\lambda_{1}\lambda_{3},\lambda_{1}\lambda_{2})$
since $\boldsymbol{\lambda}$ and $\boldsymbol{\check{\lambda}}$ as well as their permutations are connected, which can be derived from Theorem 1. Nevertheless, this criterion is still only sufficient but not necessary and rules out some frame-indifferent, isotropic polyconvex functions as numerical examples in Section 7.3 will show.
@@ 3.3.2 The incompressible case
The derivation of the above criterion for the incompressible case is also straightforward and reads:
@@ 3.4.1 The compressible case
A criterion formulated in matrix invariants can be directly motivated from Ball’s criterion by using coupled combinations of the principal stretches. Within this work, the matrix invariants ($I_{1}^{\boldsymbol{U}}=\lambda_{1}+\lambda_{2}+\lambda_{3}$, $I_{2}^{\boldsymbol{U}}=\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}+\lambda_{1}\lambda_{2},I_{3}^{\boldsymbol{U}}=J=\lambda_{1}\lambda_{2}\lambda_{3}$) of the right stretch tensor $\boldsymbol{U}=\sqrt{\boldsymbol{C}}=\sqrt{\boldsymbol{F^{T}}\cdot\boldsymbol{F}}$ are used; see, for example, [40] for more details. Let $\Psi$ be given by
@@
Certainly, the invariants of the right Cauchy–Green tensor $\boldsymbol{C}$ could also be used. However, it was shown in [16] that this criterion is not capable of approximating, e.g., energy functions exhibiting linear growth with respect to the stretch. Such linear energy growth corresponds to stress plateaus, which are observed, e.g., for phase transitions, polymers and biological materials such as the periodontal ligament [37, 24].
@@ 3.4.2 The incompressible case
The criterion reduces to the arguments $I_{1}^{\boldsymbol{U}}=\lambda_{1}+\lambda_{2}+\lambda_{3}$ and $I_{2}^{\boldsymbol{U}}=\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}+\lambda_{1}\lambda_{2}$ for incompressibility. Let $\Psi^{\textnormal{inc}}$ be given by
@@ 4 Isotropic polyconvex neural networks for incompressibility
where $\boldsymbol{x}$ is the input vector, $\boldsymbol{z}_{i}$ are the layer activations ($\boldsymbol{z}_{0}\equiv 0,$ $\boldsymbol{W}_{0}^{(\boldsymbol{z})}\equiv 0$), $\theta=\{\boldsymbol{W}_{0:k-1}^{(\boldsymbol{x})},\boldsymbol{W}_{1:k-1}^{(\boldsymbol{z})},\boldsymbol{b}_{0:k-1}\}$ are the weights, $g_{i}$ are non-linear activation functions and $k$ is the number of layers. Given this architecture, $\mathcal{NN}$ is convex in $\boldsymbol{x}$ if all $\boldsymbol{W}_{1:k-1}^{(\boldsymbol{z})}$ are non-negative and all activation functions $g_{i}$ are convex and non-decreasing [4]. ICNNs are employed in this work because of their representational power. A universal approximation theorem for convex functions was proven by Huang et al. [20] and Chen et al. [12] when ReLU or Softplus activation functions are used (see also Lemma 6). Of course, other neural network architectures ensuring convexity in the input arguments could be employed as long as the universal approximation property can be proven, see Remark 1.
@@ 4.1 Convex Signed Singular Value Neural Networks
To approximate any frame-indifferent, isotropic polyconvex function under the incompressibility constraint, the same CSSV-NN framwork as in [16], but reduced by the determinant, is proposed and hence labeled as inc-CSSV-NN. This corresponds to the polyconvexity criterion given by Corollary 2. The modified inc-CSSV-NN framework is summarized briefly below.
@@
See Tab. 1 for all combinations of $\boldsymbol{x}^{(j)}$. It should be emphasized that the same $\mathcal{NN}$ and therefore the same weights have to be used for every input permutation in order to guarantee $\Pi_{3}$-invariance. The overall inc-CSSV-NN architecture and all following neural network architectures are sketched in Fig. 3. A universal approximation theorem for the underlying inc-CSSV-NN is proven in Section 5. Such a universal approximation theorem cannot be derived for the other neural networks since they are only based on sufficient criteria. The first Piola–Kirchhoff stress tensor defined by the network potential then reads
@@
Kolmogorov–Arnold Networks (KANs) [28] currently emerge as promising alternatives to ICNNs. They were first applied to constitutive modeling by Abdolazizi et al. [2] and afterwards extended to account for polyconvexity in [42]. While the Kolmogorov–Arnold representation theorem provides a strong theoretical background, the approximation capabilities of KANs, especially under further constraints such as convexity, and limitations are yet to be explored more deeply and universal approximation theorems (for those specific cases) are to be developed. While this states an interesting, future challenge, it is beyond the scope of this work and not necessary for the present purpose since ICNNs already guarantee universal approximation. Furthermore, the required Perm(3)(3)-invariance could also be ensured via Deep Sets [48] to eventually save some computational time. However, it has not yet been investigated whether universal approximation can be achieved for the underlying constraints of $\Pi_{3}$-invariance and convexity. The proposed averaging technique in Eq. 37 is simple, effective, does not limit the network expressiveness and is computationally efficient since the network evaluations can be performed simultaneously.
@@ 4.2 Reduced Convex Signed Singular Value Neural Networks
4.3 Neural networks based on the improved Ball criterion formulated in principal stretches
The general framework for neural network models based on Theorem 5 for polyconvexity (here referred to as Ball NN and inc-Ball NN in the incompressible case) is very similar to the one shown in the previous sections. The elementary polynomials of the principal stretches are the network input instead of the signed singular values. Then, only the respective invariance has to be changed and the non-decreasing property of Corollary 5 needs to be taken into account. The Perm(3)(3)-invariance is guaranteed by setting
@@ 4.3 Neural networks based on the improved Ball criterion formulated in principal stretches
See Tab. 3 for all combinations of $\boldsymbol{x}_{\lambda}^{(j)}$. To account for the additional non-decreasing property – in addition to $\boldsymbol{W}_{1:k-1}^{(\boldsymbol{z})}$ – also all $\boldsymbol{W}_{0:k-1}^{(\boldsymbol{x})}$ have to be non-negative. This is quite restrictive for the expressiveness of ICNNs and comes on top of the fact that the improved Ball criterion is only sufficient for polyconvexity. However and for the same reasons as the reduced inc-/CSSV-NN, the model is computationally more efficient and less susceptible to local minima during optimization than the inc-/CSSV-NN but less efficient compared to the reduced inc-/CSSV-NN.
Table 3: Input permutations of the principal stretches for the neural network employed in Eq. (43) based on the improved Ball criterion for polyconvexity.
@@ 4.4 Polyconvex neural networks based on matrix invariants
However, as for the neural network models based on the improved Ball criterion, all $\boldsymbol{W}_{0:k-1}^{(\boldsymbol{x})}$ have to be non-negative to ensure the non-decreasing property. While this model, henceforth referred to as inc-UInvar NN, is the simplest and most efficient computationally, it inherently has the least expressive power.
@@ 6.1 Network architecture
CSSV-NN 7 inc-CSSV-NN 6
Reduced CSSV-NN 3 Reduced inc-CSSV-NN 2
Ball NN 7 inc-Ball NN 6
UInvar NN 3 inc-UInvar NN 2
@@ 6.2 Sobolev training
The networks are trained in Sobolev space by strain–stress tuples
@@
with the Froebenius norm $\|\bullet\|$ as it is nowadays common practice in constitutive modeling. This formulation measures deviations in stress space, thus offering a natural physical interpretation. While alternative loss definitions are possible (e.g., based on strain–energy data or a mixture of strain–energy and strain–stress), training exclusively on stress data produced the most reliable outcomes and aligns well with the availability of experimental measurements. Note that the inc-CSSV-NN framework can also be applied to approximate polyconvex hulls in the same manner as described in [16].
@@ 7 Numerical examples
In the following, several numerical examples are conducted to analyze the expressiveness, accuracy, simplicity and efficiency of the proposed models. In a first step, the models are trained on synthetic data from four well-established, classical models. Then, they are trained on Treloar’s experimental data [43] to highlight their practical applicability. Finally, an explicitly constructed energy function is provided that cannot be approximated by neural networks based on the improved Ball criterion for polyconvexity.
@@ 7.1 Classical models
leading to 52 data points (with uniform distribution within each load case). $F_{33}$ follows from incompressibility and the Lagrange multiplier $p$ from the constraint $P_{33}=0$. A separate evaluation with validation and test data sets is omitted for the noise-free synthetic data base that is directly extracted from the ground truth reference. This does not limit any of the following statements. For each combination of material model, neural network model (inc-CSSV-NN, reduced inc-CSSV-NN, inc-Ball NN and inc-UInvar NN) and architecture (1–8), 30 instances with different random weight initializations were trained using SciPy’s minimize function, which performs a constrained L-BFGS-B optimization. The number of maximum iterations was set to 1000. However, nearly every optimization stopped earlier based on SciPy’s default criteria.
@@ 7.2 Treloar data
Again, the inc-CSSV-NN, the reduced inc-CSSV-NN and the inc-Ball NN fit the material behavior perfectly, while the inc-UInvar NN fails to do so, cf. Fig. 8. This is also evident in the mean squared errors shown in Tab. 7. It should be mentioned, that the first three neural network models perform magnitudes better than all 14 classical models investigated in [41] highlighting the superior expressiveness and accuracy of neural network models. Unfortunately, an exact error comparison is not possible because the stress components included in the error [41] are not explicitly accessible. Interestingly, the test set mean squared errors are lower than the training and validation errors. This can be justified by the data distribution: the randomly selected, few test data points lie all in low stretch and stress regimes, which inherently show lower squared errors. In these regimes even the inc-UInvar NN provides a really good approximation.
So far, no significant difference between the inc-CSSV-NN, the reduced inc-CSSV-NN and the inc-Ball NN regarding the accuracy and expressiveness can be identified. For this reason, a final example is examined.
@@ 7.3 Mielke energy
and showed that it is polyconvex, despite being non-convex in $\boldsymbol{\nu}$ and, in particular, not monotone in any $\nu_{i}$. Especially due to the latter, it seems reasonable to assume that this energy cannot be approximated by neural networks that are based on the improved Ball criterion for polyconvexity, which is investigated in this section. Since the energy (54) is not differentiable, a smooth, polyconvex approximation given by
@@ 7.3.1 Incompressible Mielke energy
This time, only the inc-CSSV-NN and the reduced inc-CSSV-NN are able to reproduce the material model perfectly, cf. Fig. 9 and Tab. 6. Their lowest mean squared errors are 9.73e-05 for the inc-CSSV-NN and 6.63e-05 for the reduced version. In contrast, the lowest mean squared error achieved by the inc-Ball NNs (6.95e-03) is 71 times higher than the lowest inc-CSSV-NN error and 104 times higher than the lowest reduced inc-CSSV-NN error. The inc-Ball NN is, e.g., not able to predict the stress component $P_{11}$ for compressive strains in $11$-direction while stretched in $22$-direction, see Fig. 9 right, or decreasing stresses in the pure shear case ($\lambda_{2}=1$, Fig. 9 middle and right) accurately. This behavior is expected from the respective polyconvexity criterion. The inc-UInvar NN performs even worse (lowest mean squared error 7.44e-02) and shows only a smooth, but far off approximation of the stresses (Fig. 9). The details of the material behavior prescribed by Eq. (56) are not captured.
Finally, it can be concluded from this example that the expressivity of the inc-Ball NN is indeed limited. This indicates that the improved version of Ball’s polyconvexity criterion and thus also Ball’s original criterion is not necessary in the incompressible case. In contrast, the reduced inc-CSSV-NN approximated all incompressible energies perfectly. It hence does not contradict the ?-marked relation in Fig. 2 regarding the necessity of the reduced criterion based on the elementary symmetric polynomials of the signed singular values in the incompressible case. However, further mathematical investigations are required and rigorous proofs of these statements remain an open challenge.